Oxigen atoms in 4.90 mol in Al2O3

1
Expert's answer
2021-04-21T03:55:04-0400

1 mole Al2O3 contains 2 moles of Al and 3 moles of O.

= 3/5 × 4.90

= 2.94 moles

1 mole = 6.02 × 1023

= 2.94 × 6.02 × 1023

= 1.7699 × 1024 atoms